OpenStudy (anonymous):

Can somebody help me with 3ab^2 - 3abc factorizing it

OpenStudy (saifoo.khan):

3ab(b-c)

OpenStudy (anonymous):

ok good you know what you are doing haha i need help on a few if you would?

OpenStudy (saifoo.khan):

umm, u are doing a test or somthing? lol. i can try for sure.

OpenStudy (anonymous):

no end of term homework b^4 - ab^3 + b^2c

OpenStudy (saifoo.khan):

\[b^2 (-a b+b^2+c) \]

OpenStudy (anonymous):

could you explain how you do it?

OpenStudy (anonymous):

ok sorry that was an easier one now can you help where it is complicated? 3yz - 6z - 9 and would it be 3( y^2- 2z - 3)?

OpenStudy (saifoo.khan):

correct.

OpenStudy (anonymous):

y^3z^2 - y^2Z^2 - y^2Z^2 1= y^2z(yz - z -z^2)?

OpenStudy (lilg132):

you have to look at which of the variables appears in all of them so for example in the last one saifoo answered b appeared the most but as you can see all of them have atleast more then b^2 which means we can write b^2 outside the bracket and we need to put inside that bracket something * b^2 which will give us b^4 - ab^3 + b^2c so you would put b^2(b^2 - ab + c) when you expand it you get b^2 * b^2 = b^4 b^2 * -ab = -ab^3 b^2 * + c = b^2c that is how you check if it is correct but the main thing you have to look out for is what appears in all of them e.g. what variable or what number can be used to multiply with another variable or number to give you that equation

OpenStudy (anonymous):

Ok thank you so much... was that last one I did correct?

OpenStudy (lilg132):

yes it is correct